OpenStudy (bluescar):

A rectangle has a width of 7.9 ft and a perimeter of 27.2 ft. Enter the length of the rectangle, in decimal form, in the box. ft

OpenStudy (phi):

do you know the formula for the perimeter of a rectangle ?

OpenStudy (bluescar):

nope

OpenStudy (phi):

can you look it up and post it here ?

OpenStudy (bluescar):

p=2(1+w)

OpenStudy (phi):

you mean p= 2(L+w) (not 1) next, you can replace the letters with numbers they tell you what "p" is (p stands for perimeter) and they told you W can you fill in the numbers ?

OpenStudy (bluescar):

p=2(l+7.9)

OpenStudy (phi):

yes , but you can also replace p (perimeter) what do you get ?

OpenStudy (phi):

27.2= 2(L+7.9) now divide both sides by 2 on the left side you get 27.2 on the right side 2/2 is 1 (so we can ignore it) we get 13.6= L+7.9 last step: add -7.9 to both sides

OpenStudy (bluescar):

ok you get 27.2=2(1+7.9)

OpenStudy (bluescar):

oh

OpenStudy (phi):

yes

OpenStudy (bluescar):

won't let me paste my answer XD

OpenStudy (bluescar):

13.6+-7.9=5.7

OpenStudy (phi):

yes, OS is a pain today. I have to refresh the page but it looks like this 13.6= L+7.9 13.6 - 7.9= L+7.9-7.9 notice we keep the L and just "add on" -7.9 to both sides now simplify. we know 7.9-7.9 is 0 on the right side

OpenStudy (bluescar):

ok

OpenStudy (phi):

and as you posted , after simplifying 5.7= L or L= 5.7 the answer!

OpenStudy (bluescar):

thank you, can you help me with one more and check some of my other questions?

OpenStudy (phi):

one more. can you make a new post please?

OpenStudy (bluescar):

ok